OpenStudy (anonymous):

Solve for y in the equation –11y = –143. A. –13 B. –12 C. 12 D. 13

OpenStudy (paki):

y= 143/11....

OpenStudy (anonymous):

13, 13*111=-143

OpenStudy (warriorz13):

d.13

OpenStudy (anonymous):

*-11

OpenStudy (paki):

y= 143/11 = 13... so option D...

OpenStudy (anonymous):

so i divide 143&11?

OpenStudy (hanner_b_nanner):

D: 13!! :)

OpenStudy (anonymous):

yes

OpenStudy (warriorz13):

No do 13*11